Need some help doing and writing this project in python.

The purpose of this assignment is to assess your ability to

  • Evaluate properties of modular arithmetic relations
  • Utilize modular arithmetic to encrypt and decrypt plain text

Research the Vigenere cipher. Write a program that contains an encode and a decode function. The encode function should take a plaintext message and a key. It should return the encoded message. The decode function should take the encoded message and the key and return the decoded message

Need to be done in Python. Thanks  
Answer #1

Solution:

Vigenere Cipher:

This is a method of encryption of alphabetic strings of data.

We use a table of alphabets stored as shown below.

The table has alphabets in it stored in 26 rows differently in each row in which each alphabet is shifted
in each row.

Encryption:

The first letter of plain text is paired with the first letter of key.

We see row and column from table using first letter of plain text as row and first letter of key as column.We use this letter for encrypted text of first letter.

This process repeates for all letters in plain text.

Decryption:

We use the row in the table that corresponds to the key.

The process of decryption is as follows:

We consider the first letter from key(Use it as row value).

We consider the first letter from cipher text and find it in that row.Use that column value as plain text.

We find decryption for that letter.In similar manner we find all the decryptions for all letters.

Another implementation:

In order to eaily implement above cipher method, we convert letters into values ranging 0 to 25.

Encryption:
E=(P+K)mod 26

Where E is encryption
   P is plain text,
   K is key

Decryption:
   D=(E-K+26) mod 26

Program:

def key(s,k): #function to generate key of length equal to text
k=list(k)
if(len(s)==len(k)): #If equal length return k
return k
else:
for i in range(len(s)-len(k)): #If not equal, then add key in cyclic manner to generate it of length of s
k.append(k[i%len(k)])
return k

def encode(s,k): #method to encode
s=s.lower() #convert text to lower case
k=key(s,k) #generate key of needed length
str="abcdefghijklmnopqrstuvwxyz"
res=[]
s=list(s)
for i in range(len(s)): #For every character in list
if s[i]==" ":
res.append(" ")
continue #If space, add to result
x=(str.index(s[i])+str.index(k[i]))%26 #get the encrypted letter and append to list res
res.append(str[x])

res=''.join(res) #convert to string and return res
return res

def decode(s,k): #method to decode the text
s=encode(s,k) #encode the string
s=s.lower()
k=key(s,k) #generate key of needed length
str="abcdefghijklmnopqrstuvwxyz"
res=[]
s=list(s)
for i in range(len(s)): #For every character in s,decode it
if s[i]==" ":
res.append(" ")  #If space, add to result
continue
x=(str.index(s[i])-str.index(k[i])+26)%26 #If not space,get decrypted character and add it to list
res.append(str[x])

res=''.join(res) #convert to string and return it
return res

print(decode("HomeworkLib","code")) #call decode to decrypt text
print(encode("HomeworkLib","code"))    #call encode to encrypt text

Screenshots:

The Screenshots are attached below for reference.

Please follow them proper indentation.